OpenStudy (anonymous):

lim[Δx:0] ((2(x+Δx)^2+5Δx-2x^2)/(1(Δx))) find the limit plz help!

OpenStudy (anonymous):

\[\lim_{x\rightarrow0} \frac{(2(x+\delta x)^2}{1(\delta x)}\] Is this what you have been asked to find?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

{lim x-0] 2(x+deltax)2+5x^2-2x^2 all over delta x

OpenStudy (anonymous):

If so, I would rewrite it like this: \[\lim_{x\rightarrow0} \frac{2(x+h)^2}{h}\]

OpenStudy (anonymous):

\[\lim_{x\rightarrow0}\frac{2(x+h)^2+5x^2-2x^2}{h}\]

OpenStudy (anonymous):

limit is Δx

OpenStudy (anonymous):

not just x-0

OpenStudy (anonymous):

foil: 2(x+deltax)^2

OpenStudy (anonymous):

= 2x^2 ...+

OpenStudy (anonymous):

... 4*x*delta x +...

OpenStudy (anonymous):

big finish... what's the last term in the expansion?

OpenStudy (swag):

?